### v1

# Orientation attractors in velocity gradient driven processes for large plastic deformations of crystals

Abstract
[…]
1 Introduction
Metallic materials used in structural applications are commonly produced as polycrystals, composed of numerous grains, each with its own unique crystallographic orientation. This results in varying degrees of plastic deformation across the grains, influenced by factors such as each grain’s orientation, geometry, interaction with neighboring crystals, and the applied loading conditions [1]. Beyond this intergranular deformation heterogeneity, intragranular heterogeneity also emerges during plastic deformation, manifesting as localized, continuous, or discontinuous orientation spreads within individual grains [2].
The stability of crystallographic orientations plays a pivotal role in the evolution of material textures, particularly during plastic deformation processes. Understanding this stability is essential for characterizing how materials respond to deformation, especially in the context of both mono-and polycrystalline structures. During deformation, lattice orientations tend to stabilize around specific configurations, often referred to as lattice orientation attractors, or preferred or ideal orientations, which are crucial for predicting material behavior under stress.
Deformation-induced texture evolution arises primarily through mechanisms like grain rotation, driven by the activity of slip systems or twinning. This dynamic process leads to the formation of intragranular substructures, where new orientations deviate from the parent grain’s initial orientation. Over deformation, these substructures tend to evolve toward preferred orientations shaped by the imposed deformation, which subsequently influences the material’s mechanical properties and overall texture [3, 4, 5, 6, 7].
Crystallographic orientation is a key determinant of whether a grain will undergo uniform (homogeneous) or localized (heterogeneous) deformation. Stable orientations typically lead to more uniform deformation, whereas unstable orientations are often linked to strain localization and deformation inhomogeneity. Thus, the stability of crystallographic orientations becomes central to understanding texture development during plastic deformation.
The first steps in the investigation of texture evolution and preferred orientations were done in [8, 9, 10, 11] for a simplified planar polycrystal with two systems where the partition of shearing among the slip systems is kinematically determined. Other papers provides insight into how initial crystallographic orientations affect the stability of lattice configurations during deformation for a specific loading as equibiaxial tension [12, 13, 14, 15], equal-channel angular extrusion channel [16, 17] or torsion [18, 19]. Generally these simulations use an Eulerian framework with a null rotation rate to isolate the effects of external forces on orientation evolution. An important effort was also done for modeling texture evolution due to dynamic recrystallization of ice to deduce the existence of orientation attractors that maximizes the resolved shear stress on the easiest slip system in the crystal, see [20, 21, 22, 23].
Conventionally, large deformation simulations are performed in a Lagrangian framework, where the mesh follows the deforming material. While effective for moderate strains, as deformation progresses, mesh distortion introduces significant errors and ultimately leads to simulation failure. To address these challenges, this paper adopts an Eulerian approach, which is particularly effective for simulating extreme deformations. In the Eulerian formulation, a fixed grid allows materials to flow through without distorting the mesh, making it ideal for analyzing complex material behavior during severe plastic deformation.
The elastic-(visco-)plastic model needs a permanent interplay between both Eulerian (more appropriate for plasticity) and Lagrangian (more appropriate for elasticity) descriptions. However, in applications involving large deformations of metals the elastic component of the deformation is small with respect to the inelastic one, and can be neglected by using a rigid-(visco)plastic approach (see for instance [24, 25, 26]). This simplified model takes important theoretical and numerical advantages by using only one (Eulerian) configuration. The present work uses the Eulerian framework in modeling large deformations in crystal plasticity introduced by [27] and uses the high-resolution numerical methods developed in [28].
The aim of this paper is to investigate the stability of lattice orientation in velocity gradient-driven processes for very large deformations. Our objective is to provide a general framework for the stability analysis of crystal lattices that can model the evolution toward stable (ideal or preferred) orientations. To simplify our analysis, only perfectly (visco-)plastic models, without hardening/softening effects, will be considered. Since the analytical stability analysis of a 3D crystal is too complex, a more detailed investigation was conducted for a simplified 2D model with only three slip systems. This model is simple enough to permit analytical analysis yet rich enough to describe the principal features of a 3D crystal. We characterize the stable stationary orientations and deduce their basins of attraction, which are essential in predicting the final texture of poly/monocrystals.
Let us outline the paper’s content. In Section 2, we recall the 3D rigid-viscoplastic crystal model from [27] and its 2D simplification, omitting hardening and softening effects. A key contribution is our analytic decomposition of strain rate into slip rates for the 2D model. Section 3 presents the stability analysis framework for lattice orientation in velocity gradient-driven processes. Since the general 3D case is too complex for analytic treatment, Section 4 provides detailed stability analysis only for the simplified 2D model, characterizing stable orientations and their attraction basins. Section 5 presents numerical simulations illustrating our theory through three cases: (1) a polycrystal with homogeneous velocity gradient but heterogeneous initial orientations, (2) a monocrystal with void growth under radial loading, and (3) slip band formation in a single crystal under uniaxial loading.
2 Eulerian rigid-plastic approach of crystal plasticity
Consider a single crystal at time $t=0$, free of any surface tractions and body forces and let choose this configuration, say ${\mathcal{K}}_{0}\subset\mathbb{R}^{3}$ as reference configuration of the crystal. Let ${\mathcal{K}}={\mathcal{K}}(t)\subset\mathbb{R}^{3}$ denote the current configuration. The incorporation of lattice features is achieved through a multiplicative decomposition of the total deformation gradient ${\boldsymbol{F}}$ into elastic and plastic components:
${\boldsymbol{F}}={\boldsymbol{F}}^{e}{\boldsymbol{P}}.$ (2.1)
This decomposition implies a two-stage deformation process. First, ${\boldsymbol{P}}$ transforms the initial reference state ${\mathcal{K}}_{0}$ to an intermediate state $\tilde{{\mathcal{K}}}$, characterized by plastic deformation only with no change in volume. ${\boldsymbol{P}}$ is called the (visco)plastic deformation with respect to the reference configuration of a material neighborhood of the material point $X$ at time $t$. Then, ${\boldsymbol{F}}^{e}$ brings the body to the final configuration ${\mathcal{K}}$ through elastic deformation and rigid lattice rotation, i.e. ${\boldsymbol{F}}^{e}={\boldsymbol{R}}{\boldsymbol{U}}^{e}$ where ${\boldsymbol{R}}$ denotes the rotation of the crystal axes with respect to its isoclinic orientation.
Following [29, 30], ${\boldsymbol{P}}$ is assumed to leave the underlying lattice structure undeformed and unrotated, ensuring the uniqueness of the decomposition in (2.1). The unique feature of CP theory is its construction of the plastic component ${\boldsymbol{P}}$ by constraining dislocation kinematics. Plastic flow evolves along pre-selected slip directions via volume-preserving shears, leaving the crystal lattice undistorted and stress-free [31].
Since in applications involving large deformations, the elastic component of the deformation is small with respect to the inelastic one, it can be neglected and a rigid-viscoplastic approach will be adopted (such a hypothesis is generally used e.g. Hutchinson [24], Lebensohn and Tome [25], Kok et al. [26], etc.). That means that we neglect the elastic lattice strain ${\boldsymbol{U}}^{e}$ by supposing that ${\boldsymbol{U}}^{e}\approx{\boldsymbol{I}}$. This leads to the following decomposition for the deformation gradient ${\boldsymbol{F}}$ (see for example Kok et al. [26]):
${\boldsymbol{F}}={\boldsymbol{R}}{\boldsymbol{P}}.$ (2.2)
Such a hypothesis is valid, as during forming or other industrial processes, the elastic component of deformation is negligibly small (typically on the order of 10-3) compared to the plastic component (typically greater than 10-1). It should also be noted that, once the elastic–plastic transition is complete, the stress evolution within the grains is governed by plastic relaxation (see Tomé and Lebensohn [25]).
2.1 Eulerian description of the lattice rotations
Crystal slip systems are labeled by integers $s=1,...,N$, with $N$ denoting the number of slip systems. Each slip system $s$ is specified by the unit vectors $({\boldsymbol{b}}_{s}^{0},{\boldsymbol{m}}_{s}^{0})$, where ${\boldsymbol{b}}_{s}^{0}$ is in the slip direction and ${\boldsymbol{m}}_{s}^{0}$ is normal to the slip plane in the perfect undeformed lattice. Since the viscoplastic deformation does not produce distortion or rotation of the lattice, the mean lattice orientation is the same in the reference and intermediate configurations and is specified by $({\boldsymbol{b}}_{s}^{0},{\boldsymbol{m}}_{s}^{0}),s=1...N$.
Let ${\boldsymbol{b}}_{s}={\boldsymbol{b}}_{s}(t)$ and ${\boldsymbol{m}}_{s}={\boldsymbol{m}}_{s}(t)$, the glide direction and glide plane normal, respectively in the deformed configuration i.e. ${\boldsymbol{b}}_{s}(0)={\boldsymbol{b}}_{s}^{0}$ and ${\boldsymbol{m}}_{s}(t)={\boldsymbol{m}}_{s}^{0}$ Since elastic effects are neglected,
${\boldsymbol{b}}_{s}={\boldsymbol{R}}{\boldsymbol{b}}_{s}^{0},\quad{\boldsymbol{m}}_{s}={\boldsymbol{R}}{\boldsymbol{m}}_{s}^{0},$ (2.3)
Note that according to (2.3), ${\boldsymbol{b}}_{s}$ and ${\boldsymbol{m}}_{s}$ are unit vectors. Furthermore,
${\boldsymbol{b}}_{s}\otimes{\boldsymbol{m}}_{s}={\boldsymbol{R}}\left({\boldsymbol{b}}_{s}^{0}\otimes{\boldsymbol{m}}_{s}^{0}\right){\boldsymbol{R}}^{T}.$ (2.4)
We seek to express the lattice evolution equations only in terms of vector and tensor fields associated with the current configuration. Let ${\boldsymbol{v}}={\boldsymbol{v}}(t,x)$, the Eulerian velocity field, ${\boldsymbol{L}}$ the velocity gradient, ${\boldsymbol{D}}$ the rate of deformation, and ${\boldsymbol{W}}$ the spin tensor
${\boldsymbol{L}}={\boldsymbol{L}}({\boldsymbol{v}})=\nabla{\boldsymbol{v}},\quad{\boldsymbol{D}}={\boldsymbol{D}}({\boldsymbol{v}})=(\nabla{\boldsymbol{v}})^{symm},\quad{\boldsymbol{W}}={\boldsymbol{W}}({\boldsymbol{v}})=(\nabla{\boldsymbol{v}})^{skew}.$ (2.5)
The viscoplastic deformation is due to slip only; the slip contribution to the viscoplastic deformation being ([32], [33])
$\dot{{\boldsymbol{P}}}{\boldsymbol{P}}^{-1}=\sum_{s=1}^{N}\dot{\gamma}^{s}{\boldsymbol{b}}_{s}^{0}\otimes{\boldsymbol{m}}_{s}^{0},$ (2.6)
where $\dot{\gamma}^{s}=\dot{\gamma}^{s}(t)$ is the viscoplastic shear rate on the slip system $s$. If we denote by
${\boldsymbol{M}}_{s}=\left({\boldsymbol{b}}_{s}\otimes{\boldsymbol{m}}_{s}\right)^{symm},\quad{\boldsymbol{Q}}_{s}=\left({\boldsymbol{b}}_{s}\otimes{\boldsymbol{m}}_{s}\right)^{skew}.$ (2.7)
then, using ${\boldsymbol{L}}=\dot{{\boldsymbol{F}}}{\boldsymbol{F}}^{-1}=\dot{{\boldsymbol{R}}}\;{\boldsymbol{R}}^{T}+{\boldsymbol{R}}\dot{{\boldsymbol{P}}}{\boldsymbol{P}}^{-1}\;{\boldsymbol{R}}^{T}$, Eqs. (2.4) and (2.6), the rate of deformation ${\boldsymbol{D}}$ can be written as
${\boldsymbol{D}}=\sum_{s=1}^{N}\dot{\gamma}^{s}{\boldsymbol{M}}_{s}.$ (2.8)
Taking the anti-symmetric part of ${\boldsymbol{L}}$, we obtain that the spin tensor is ${\boldsymbol{W}}=\dot{{\boldsymbol{R}}}\;{\boldsymbol{R}}^{T}+\sum_{s=1}^{N}\dot{\gamma}^{s}{\boldsymbol{Q}}_{s}$ and a differential equation for the rotation tensor ${\boldsymbol{R}}$:
$\dot{{\boldsymbol{R}}}=({\boldsymbol{W}}-\sum_{s=1}^{N}\dot{\gamma}^{s}{\boldsymbol{Q}}_{s}){\boldsymbol{R}}.$ (2.9)
The evolution equations (2.9) describe the evolution of the lattice in terms of vector and tensor fields associated with the current configuration.
2.2 Plastic and visco-plastic flow rules
In order to complete the model, we need to provide the constitutive equation for the slip rate $\dot{\gamma_{s}}$ as a function of $\tau_{s}$, the stress component acting on the slip plane of normal ${\boldsymbol{m}}_{s}$ in the slip direction ${\boldsymbol{b}}_{s}$. In the current configuration, $\tau_{s}$ is expressed as
$\tau^{s}={\boldsymbol{\sigma}}:{\boldsymbol{M}}_{s},$ (2.10)
where ${\boldsymbol{\sigma}}={\boldsymbol{\sigma}}(t)$ is the Cauchy stress tensor acting in the current configuration ${\mathcal{K}}$ while ${\boldsymbol{M}}_{s}$ is defined by (2.7). Note that $\{\tau^{s}\}_{s=\overline{1,N}}$ are not independent; they belong to a fifth dimensional space of $\mathbb{R}^{N}$ corresponding to the dimension of the space of deviatoric stresses.
To determine the shear strain rates $\dot{\gamma}_{\alpha}$ relative to the local stress, a constitutive law is needed. Various proposals exist, ranging from phenomenological to more physically based approaches. One simple phenomenological approach assumes that $\dot{\gamma}_{\alpha}$ depends on the stress only through the resolved shear stress $\tau^{s}$.
In rigid-plastic formulations, it is assumed that the onset of plastic flow of a slip system $s$ is governed by Schmid law: the slip system $s$ is active if and only if $|\tau^{s}|=\tau_{c}^{s}$, i.e.
$\dot{\gamma_{s}}(|\tau^{s}|-\tau_{c}^{s})=0,\quad\dot{\gamma_{s}}\tau^{s}\geq 0,\quad|\tau^{s}|-\tau_{c}^{s}\leq 0,$ (2.11)
where $\tau_{c}^{s}$ is the slip resistance (also called critical resolved shear stress or CRSS). For a given time $t$, the $\tau_{c}^{s}$ are material constants. Thus, the planes $|\tau^{s}|=\tau_{c}^{s}$ are the facets of the current yield surface of the single crystal in the stress space.
Since the resolved shear stresses $\tau^{s}$ are not independent, the shear rates $\dot{\gamma^{s}}$ given by the viscoplastic flow rule (2.11) are not independent; they have to satisfy the kinematic constraint (2.8). Given the rate of deformation ${\boldsymbol{D}}$, the slip rates $\dot{\gamma^{s}}$ can be determined by minimizing the internal plastic dissipation power
$J_{Schmid}(\dot{\gamma}^{1},\dot{\gamma}^{2},...,\dot{\gamma}^{N})=\sum_{s=1}^{N}\tau_{c}^{s}|\dot{\gamma}^{s}|,$ (2.12)
over $\mathbb{R}^{N}$ under the constraint (2.8). The above functional is neither strongly convex or differentiable and the solution could not be unique. Additional assumptions are needed in order to restrict the number of solutions.
One way to overcome this difficulty of determining the active slip systems problem is to adopt a rate-dependent approach for the constitutive response of the single crystal. A widely used rate-dependent (viscoplastic) model is the Norton type model, which relates the shear strain rate $\dot{\gamma}^{s}$ on a slip system $s$ to the resolved shear stress $\tau^{s}$ through a power-law (see Asaro and Needleman [34])
$\dot{\gamma}^{s}=\dot{\gamma}_{0}^{s}\,{\left|\frac{\tau^{s}}{\tau_{c}^{s}}\right|}^{n}\,\mbox{sign}(\tau^{s}),$ (2.13)
where $\dot{\gamma}_{0}$ is a reference shear strain rate, while the exponent n has a fixed value. Since the internal plastic dissipation power reads
$J_{Norton}(\dot{\gamma}^{1},\dot{\gamma}^{2},...,\dot{\gamma}^{N})=\frac{n}{n+1}\sum_{s=1}^{N}\tau_{c}^{s}|\dot{\gamma}^{s}|\left|\frac{\dot{\gamma}^{s}}{\dot{\gamma}_{0}^{s}}\right|^{\frac{1}{n}},$ (2.14)
we remark that $J_{Norton}\to J_{Schmid}$ for large values of $n$ ($n\to\infty$) which means that the Norton law is a visoplastic regularization of the Schmidt law. Note that the plastic dissipation functional is strongly convex and differentiable which means the slip rates are the unique solution of the minimization problem of the plastic dissipation power $J_{Norton}$ over $\mathbb{R}^{N}$ under the constraint (2.8).
Another regularization of the Schmid law can be done by using a Perzyna-like viscoplastic law of the form.
$\dot{\gamma}^{s}=\dfrac{1}{\eta_{s}}\left[|\tau^{s}|-\tau_{c}^{s}\right]_{+}\mbox{sign}(\tau^{s}),$ (2.15)
where $\eta_{s}$ is the viscosity, which may depend on the slip rate, and $[\;x\;]_{+}=(x+|x|)/2\;$ denotes the positive part of any real number $x$. Note that the viscoplastic flow rule (2.15) is the visco-plastic extension of the rigid-plastic Schmid law using an overstress approach. The physical motivation for the dependence of the viscoplastic shear rate on the overstress $(\tau^{s}-\tau_{c}^{s})$ was provided by Teodosiu and Sidoroff [33] based on an analysis of the microdynamics of crystals defects.
The internal power is given by
$J_{Perzyna}(\dot{\gamma}^{1},\dot{\gamma}^{2},...,\dot{\gamma}^{N})=\sum_{s=1}^{N}\frac{\eta_{s}}{2}|\dot{\gamma}^{s}|^{2}+\tau_{c}^{s}|\dot{\gamma}^{s}|,$ (2.16)
and we remark that $J_{Perzyna}\to J_{Schmid}$ for a vanishing viscosity $\eta_{s}$ ($\eta\to 0$). The above functional is strongly convex, hence the slip rates are the unique solution of the minimization problem of the plastic dissipation power $J_{Perzyna}$ over $\mathbb{R}^{N}$ under the constraint (2.8).
Finally we deduce that the slip rates $\dot{\gamma}^{1},\dot{\gamma}^{2},...,\dot{\gamma}^{N}$ can be obtained by an optimization problem involving the plastic dissipation power
$(\dot{\gamma}^{1},\dot{\gamma}^{2},...,\dot{\gamma}^{N})=\arg\min_{{\boldsymbol{D}}=\sum_{s=1}^{N}g^{s}{\boldsymbol{M}}_{s}}J(g^{1},g^{2},...,g^{N}).,$ (2.17)
where $J$ is one of the plastic dissipation power functionals $J_{Schmid},J_{Norton}$ or $J_{Perzyna}$.
2.3 The 2-D model with 3 slip systems
The analysis of 3D crystalline systems with $N$ slip systems presents significant theoretical and physical challenges. To facilitate understanding of the complex deformation phenomena, simplifications are often beneficial. In this work, we introduce a two-dimensional model featuring $N=3$ slip systems. Unlike the planar crystal model with $N=2$ slip systems proposed by Prantil et al. [8], the 3-slip system model, used here, is not kinematically determined—meaning the distribution of shear across the slip systems is not predetermined.
Let us denote with $\phi$ and $-\phi$ the angles formed by slip system $r=1$ with the other two systems $r=2,3$ and let $\theta$ be the angle formed by the slip system $1$ with the $Ox_{1}$ axis, see Fig. 1a. The three composite in-plane slip systems ${\boldsymbol{b}}_{1},{\boldsymbol{b}}_{2},{\boldsymbol{b}}_{3}$ are specified by the angles $\theta,\theta+\phi,\theta-\phi$.
2.3.1 Lattice orientation equation
The main simplification for the 2-D problem comes from the latice rotation ${\boldsymbol{R}}$ which is now a rotation ${\boldsymbol{R}}(\theta,{\boldsymbol{e}}_{3})$ with angle $\theta$ along $0x_{3}$ axis and we have
${\boldsymbol{Q}}_{r}=\dfrac{1}{2}\left((1,0)\otimes(0,1)-(0,1)\otimes(1,0)\right).$ (2.18)
for all $r=1,2,3$ and (2.9) has a much more simpler form
$\dot{\theta}=\frac{\partial\theta}{\partial t}+{\boldsymbol{v}}\cdot\nabla\theta=\frac{1}{2}\left(\sum_{r=1}^{3}\dot{\gamma}^{r}-(\frac{\partial v_{1}}{\partial x_{2}}-\frac{\partial v_{2}}{\partial x_{1}})\right).$ (2.19)
2.3.2 Specific applications
We have in mind two situations where the above model is physically relevant: (i) in-plane deformation of an FCC crystal, and (ii) slip in the basal plane of a hexagonal crystal. We first examine the case of in-plane deformation in FCC crystals, and subsequently consider the second scenario.
In-plane deformation of FCC crystals: Rice [32] showed that certain pairs of the three-dimensional systems that are potentially active need to combine in order to achieve plane-strain deformation. For a F.C.C. crystal, with 12 potentially active slip systems, we consider $Ox_{3}$ axis to be parallel to $[110]$ in the crystal basis, which means that the plane-strain plane $(Ox_{1}x_{2})$ is the plane $[\bar{1}10]-[001]$. Some geometrical constraints (see [28]) have to are satisfied such that $N=3$ pairs of composite systems will give deformation in the plane $(Ox_{1}x_{2})$. The in-plane system $r=1$ is formed by $(k,l)=(3,12)$, $r=2$ is formed by $(k,l)=(4,5)$, and $r=3$ is formed by $(k,l)=(7,8)$ (here $k,l$ are number of slipping directions for the 3-D FCC crystal). We also suppose in what follows that the systems $1,2,6,9,10$ and $11$ are not active. For all $r=1,2,3$ we have denoted by ${\boldsymbol{b}}_{r},{\boldsymbol{m}}_{r}$ the normalized projections of the corresponding three-dimensional slip directions and normal directions $(k,l)$ onto the $x_{1}x_{2}$-plane. The angle $\phi$ between the slipping sytem $1$ and $2$ is
$\phi=\arctan(\sqrt{2})\approx 54.7^{\circ}.$ (2.20)
Since there are some scalar factors between the first two components of the in-plane systems and the 3-dimensional ones given by $q_{1}=\frac{1}{\sqrt{3}},\quad q_{2}=q_{3}=\frac{\sqrt{3}}{2},$ the 2-D composite slipping rate $\dot{\gamma}^{r}$ corresponds to the 3-D slipping rate $\dot{\gamma}^{k}$ multiplied by $2q_{r}$. As it follows from [28] the 2-D yield limit $\tau_{c}^{r}$ corresponds to the 3-D yield limit $\tau_{c}^{k}$ divided by $q_{r}$ while the 2-D viscosity $\eta^{r}$ corresponds to the 3-D viscosity $\eta^{k}$ divided by $q_{r}^{2}$.
Slip in the basal plane of a hexagonal crystal.: Alternatively, the 2D model with three slip systems corresponds to hexagonal close-packed (HCP) crystals (such as Ti, Mg, Zr, etc.) under plane strain conditions, where the deformation plane is aligned with the basal plane (0001) of the hexagonal lattice. In this setting—experimentally studied in [35]—plastic strain is primarily accommodated by the three prismatic slip systems, namely the $(10\bar{1}0)\langle 1\bar{2}10\rangle$ slip family. Each of these slip systems can operate independently of the others, as their strain contributions are consistent with plane strain conditions. It is worth noting that combinations of other HCP slip systems from the basal or pyramidal families could, in principle, also produce plane strain. However, these alternatives would require significantly more energy and are thus not observed experimentally. For this reason, such slip families are not considered in the present work. The symmetry of the prismatic slip systems leads to
$\phi=\pi/3=60^{\circ}.$ (2.21)
2.3.3 Rate of deformation slip rate decomposition
Another important simplification obtained in [28] for this model is the availability of analytical expressions for the slip rates on the individual composite systems, denoted by $\dot{\gamma}_{r}$, for a given rate of deformation ${\boldsymbol{D}}$. In the Eulerian basis ${\boldsymbol{e}}_{1},{\boldsymbol{e}}_{2}$, this rate of deformation is expressed as
${\boldsymbol{D}}=D_{11}({\boldsymbol{e}}_{1}\otimes{\boldsymbol{e}}_{1}-{\boldsymbol{e}}_{2}\otimes{\boldsymbol{e}}_{2})+D_{12}({\boldsymbol{e}}_{1}\otimes{\boldsymbol{e}}_{2}+{\boldsymbol{e}}_{2}\otimes{\boldsymbol{e}}_{1}),$ (2.22)
and a given lattice orientation $\theta$. The slip rates $\dot{\gamma}_{r}$ ($r=1,2,3$) can be determined by minimizing the (visco-)plastic power $J$ under the constraints (2.8). Let us first remark that ${\boldsymbol{M}}_{r}={\boldsymbol{R}}(\theta,{\boldsymbol{e}}_{3})\tilde{{\boldsymbol{M}}}_{r}{\boldsymbol{R}}(-\theta,{\boldsymbol{e}}_{3})$, where $\tilde{{\boldsymbol{M}}}_{r}=\left(\tilde{{\boldsymbol{b}}}_{r}\otimes\tilde{{\boldsymbol{m}}}_{r}\right)^{symm}$ and $\tilde{{\boldsymbol{b}}}_{r}$ corresponds to the orientation $\theta=0$, i.e. $\tilde{{\boldsymbol{b}}}_{1}=(1,0),\tilde{{\boldsymbol{b}}}_{3}=(\cos(\phi),\sin(\phi))$ and $\tilde{{\boldsymbol{b}}}_{2}=(\cos(\phi),-\sin(\phi))$. If the initial orientation of the lattice is given by $\theta_{0}$ then ${\boldsymbol{M}}_{r}^{0}={\boldsymbol{R}}(\theta_{0},{\boldsymbol{e}}_{3})\tilde{{\boldsymbol{M}}}_{r}{\boldsymbol{R}}(-\theta_{0},{\boldsymbol{e}}_{3})$. The strain rate decomposition (2.8) could be written as
${\boldsymbol{R}}(-\theta,{\boldsymbol{e}}_{3}){\boldsymbol{D}}{\boldsymbol{R}}(\theta,{\boldsymbol{e}}_{3})=\sum_{r=1}^{3}\dot{\gamma}^{r}\tilde{{\boldsymbol{M}}}_{r}^{0}.$ (2.23)
Let us first decompose the three dimensional vector
$\dot{\Gamma}=(\dot{\gamma}_{1},\dot{\gamma}_{2},\dot{\gamma}_{3})=s{\boldsymbol{S}}^{\phi}+c{\boldsymbol{C}}^{\phi}+z{\boldsymbol{Z}}^{\phi}$ (2.24)
in the basis $\{{\boldsymbol{S}}^{\phi},{\boldsymbol{C}}^{\phi},{\boldsymbol{Z}}^{\phi}\}$, given by ${\boldsymbol{S}}^{\phi}=(0,\sin(2\phi),-\sin(2\phi))$, ${\boldsymbol{C}}^{\phi}=(1,\cos(2\phi),\cos(2\phi))$ and ${\boldsymbol{Z}}^{\phi}={\boldsymbol{S}}^{\phi}\wedge{\boldsymbol{C}}^{\phi}=(-\sin(4\phi),\sin(2\phi),\sin(2\phi))$. Substitution of (2.24) into (2.23) leads to:
$s(\theta)=-2\frac{\cos(2\theta)D_{11}+\sin(2\theta)D_{12}}{|{\boldsymbol{S}}^{\phi}|^{2}},\quad\quad c(\theta)=2\frac{\cos(2\theta)D_{12}-\sin(2\theta)D_{11}}{|{\boldsymbol{C}}^{\phi}|^{2}},$ (2.25)
where $|{\boldsymbol{S}}^{\phi}|^{2}=2\sin^{2}(2\phi)$ and $|{\boldsymbol{C}}^{\phi}|^{2}=1+2\cos^{2}(2\phi)$.
Since $s$ and $c$ are obtained from the kinematic condition (2.23) the slip rates $\dot{\Gamma}$ from (2.24) is a function of $z$ hence the internal power $J$ is a function of $z$ only, i.e. $J(\dot{\gamma}^{1},\dot{\gamma}^{2},\dot{\gamma}^{3})=:J(z)$. For instance for the Schmid model we have
$J_{Schmid}(z)=\sum_{r=1}^{3}\tau_{c}^{r}|s(\theta)S_{r}^{\phi}+c(\theta)C_{r}^{\phi}+zZ_{r}^{\phi}|.$ (2.26)
Note that $z\to J^{Schmid}(z)$ is differentiable everywhere with the exception of three points, $z_{1}^{\theta},z_{2}^{\theta}$ and $z_{3}^{\theta}$ where $s(\theta)S_{r}^{\phi}+c(\theta)C_{r}^{\phi}+zZ_{r}^{\phi}=0$, given by
$z_{1}^{\theta}=-\frac{c(\theta)}{\sin(4\phi)},\quad z_{2}^{\theta}=\frac{s(\theta)\sin(2\phi)+c(\theta)\cos(2\phi)}{\sin(2\phi)},\quad z_{3}^{\theta}=\frac{-s(\theta)\sin(2\phi)+c(\theta)\cos(2\phi)}{\sin(2\phi)}.$
However, the left and right derivatives $J_{Schmid}^{\prime}(z_{r}^{\theta}-),J_{Schmid}^{\prime}(z_{r}^{\theta}+)$ exist and can be computed from
$J_{Schmid}^{\prime}(z)=\sum_{r=1}^{3}\tau_{c}^{r}|Z_{r}^{\phi}|\frac{z-z_{r}^{\theta}}{|z-z_{r}^{\theta}|}.$
Since $z\to J_{Schmid}(z)$ is piece-wise linear the minimum of this convex function is $z(\theta)=z_{m}^{\theta}$ where $z_{m}^{\theta}$ is such that $J^{Schmid}(z_{m}^{\theta})=\min_{r=1,2,3}J^{Schmid}(z_{r}^{\theta})$ and we get
$z(\theta)=z_{r}^{\theta}\quad\mbox{ if }\quad J_{Schmid}^{\prime}(z_{r}^{\theta}-)J_{Schmid}^{\prime}(z_{r}^{\theta}+)\leq 0.$ (2.27)
For the Perzyna viscoplastic regularization the power dissipation reads
$J_{Perzyna}(z)=\sum_{r=1}^{3}\frac{\eta_{r}}{2}|s(\theta)S_{r}^{\phi}+c(\theta)C_{r}^{\phi}+zZ_{r}^{\phi}|^{2}+\tau_{c}^{r}|s(\theta)S_{r}^{\phi}+c(\theta)C_{r}^{\phi}+zZ_{r}^{\phi}|.$ (2.28)
and the general expression of $J^{\prime}(z)$ as:
$J_{Perzyna}^{\prime}(z)=A^{\theta}+B^{\phi}z+\sum_{r=1}^{3}\tau_{c}^{r}|Z_{r}^{\phi}|\frac{z-z_{r}^{\theta}}{|z-z_{r}^{\theta}|},$
with $A^{\theta}=\sum_{r=1}^{3}\eta_{r}(s(\theta)S_{r}^{\phi}+c(\theta)C_{r}^{\phi})Z_{r}^{\phi},\;B^{\phi}=\sum_{r=1}^{3}\eta_{r}(Z_{r}^{\phi})^{2}.$ We can now solve the equation $(J^{Perzyna})^{\prime}(z)=0$ to get the analytical expression of $z(\theta)$
$z(\theta)=\left\{\begin{aligned} &z_{r}^{\theta}\quad\mbox{ if }\quad J_{Perzyna}^{\prime}(z_{r}^{\theta}-)J_{Perzyna}^{\prime}(z_{r}^{\theta}+)\leq 0,\\ &\frac{1}{B^{\phi}}\left[-A^{\theta}+\sum_{r=1}^{3}\tau_{0}^{r}|Z_{r}^{\phi}|\frac{J_{Perzyna}^{\prime}(z_{r}^{\theta}+)}{|J_{Perzyna}^{\prime}(z_{r}^{\theta}+)|}\right]\quad\hbox{otherwise }\end{aligned}\right.$ (2.29)
Further substitution of (2.25) and (2.27) (or (2.29) ) into (2.24) gives the analytical decomposition of the strain rate ${\boldsymbol{D}}({\boldsymbol{v}})$ into the three slip rates $\dot{\gamma}_{1},\dot{\gamma}_{2},\dot{\gamma}_{3}$
$(\dot{\gamma}_{1}(\theta),\dot{\gamma}_{2}(\theta),\dot{\gamma}_{3}(\theta))=s(\theta){\boldsymbol{S}}^{\phi}+c(\theta){\boldsymbol{C}}^{\phi}+z(\theta){\boldsymbol{Z}}^{\phi}$ (2.30)
according to the constitutive equation (2.11) (or (2.15)).
2.3.4 Slip rates as function of the lattice orientation
Starting from the above decomposition of the strain rate into slip rates, we investigate here the dependence of the slip rates on the lattice orientation, which will be a key tool in the stability analysis developed in Section 4.
Figure 1: (a) Two-dimensional model with three slip systems. (b) Definition of angles: $\psi$, the principal strain rate direction, and $\theta$, the lattice orientation. (c) Reference slip rates $\dot{\gamma}_{1}^{\text{ref}}(\theta)$, $\dot{\gamma}_{2}^{\text{ref}}(\theta)$, and $\dot{\gamma}_{3}^{\text{ref}}(\theta)$ as functions of $\theta$ for an FCC crystal. (d) Reference slip rate $\dot{\gamma}_{1}^{\text{ref}}(\theta)$ as a function of $\theta$ in the Perzyna model, for different values of the viscosity parameter: $\eta=0.5\eta^{\text{ref}}$ (magenta), $\eta=0.4\eta^{\text{ref}}$ (green), $\eta=0.33\eta^{\text{ref}}$ (blue), $\eta=0.30\eta^{\text{ref}}$ (red), and $\eta=0$ (i.e., the Schmidt model, yellow), with $\eta^{\text{ref}}=\tau_{c}/\dot{\Gamma}^{\text{ref}}$. (e) The function $\theta\mapsto\mathcal{G}(\theta)$ for an FCC crystal (green, $\phi=54.7^{\circ}$) and a hexagonal crystal (blue, $\phi=60^{\circ}$).
The above formula, which gives the slip rate decomposition of the strain rate given by (2.22) in the Eulerian basis, is useful for the FE computations and plays a key role in the numerical algorithm. For the stability analysis is more convenient to represent the strain rate ${\boldsymbol{D}}$ in its principal directions, denoted by ${\boldsymbol{d}}_{1},{\boldsymbol{d}}_{2}$, as
${\boldsymbol{D}}=d({\boldsymbol{d}}_{1}\otimes{\boldsymbol{d}}_{1}-{\boldsymbol{d}}_{2}\otimes{\boldsymbol{d}}_{2}),$ (2.31)
where $d>0$ is the principal rate of deformation. We denote by $\psi$ the angle between ${\boldsymbol{d}}_{1}$ and ${\boldsymbol{e}}_{1}$ (see Fig. 1(b). To compute the slip rate decomposition in this configuration les us consider first the non-dimensional case $D_{12}=0,D_{11}=1$, called in the following "reference case", corresponding to $d=1$ and $\psi=0$. Let denote by
$(\dot{\gamma}_{1}^{ref}(\theta),\dot{\gamma}_{2}^{ref}(\theta),\dot{\gamma}_{3}^{ref}(\theta))=s^{ref}(\theta){\boldsymbol{S}}^{\phi}+c^{ref}(\theta){\boldsymbol{C}}^{\phi}+z^{ref}(\theta){\boldsymbol{Z}}^{\phi},$ (2.32)
where $s^{ref},c^{ref}$ and $z^{ref}$ are given by the above formulas for $D_{12}=0,D_{11}=1$. In Fig. 1(c), we plot the reference slip rates $\dot{\gamma}_{1}^{ref}(\theta)$, $\dot{\gamma}_{2}^{ref}(\theta)$, and $\dot{\gamma}_{3}^{ref}(\theta)$ for the Schmidt model in the case of an FCC crystal. We note that for 12 orientations ($\theta=\phi/4,\pi/4,\pi/2-\phi/4,\pi/2+\phi/4,3\pi/4,\pi-\phi/4,\dots$), only one slip system is active. For any other orientation, two slip systems are active, and the maximum/minimum of the reference slip rate is 2 and -2, respectively.
Moreover, the dependence of the slip rates on the lattice orientation is not smooth everywhere: at angles where only one slip system is active, the derivative is discontinuous. This derivative discontinuity, which is specific to the Schmid rigid-plastic law, is smoothed out in the case of Perzyna regularization. To illustrate this, we have plotted in Fig. 1d the first reference slip rate $\dot{\gamma}_{1}^{ref}$ as a function of the orientation $\theta$ for different values of the viscosity $\eta$, in the case of an FCC crystal. We observe that for larger viscosity values (greater than 50% of $\tau_{c}/\dot{\Gamma}^{ref}$), the Schmid slip rate is not recovered by the Perzyna visco-plastic regularization, while for smaller viscosities (less than 25% of $\tau_{c}/\dot{\Gamma}^{ref}$), the rigid-plastic and visco-plastic models become indistinguishable.
We can now write the (dimensional) slip rate decomposition corresponding to the strain rate representation (2.31) as:
$(\dot{\gamma}_{1}(\theta),\dot{\gamma}_{2}(\theta),\dot{\gamma}_{3}(\theta))=d\left(\dot{\gamma}_{1}^{ref}(\theta-\psi),\dot{\gamma}_{2}^{ref}(\theta-\psi),\dot{\gamma}_{3}^{ref}(\theta-\psi)\right).$ (2.33)
Moreover, it is convenient (see the next section) to define ${\mathcal{G}}$—the sum of the reference slip rates, as it appears in (2.19)—given by:
${\mathcal{G}}(\theta)=\sum_{s=1}^{3}\dot{\gamma}_{s}^{ref}(\theta),$ (2.34)
In Fig. 1(e), we plot the non-dimensional function ${\mathcal{G}}$ for the in-plane deformation of an FCC crystal (green, $\phi=54.7^{\circ}$) and of a hexagonal crystal (blue, $\phi=60^{\circ}$).
2.4 Initial and boundary value problem formulation
We begin by presenting the equations governing the motion in a domain ${\mathcal{D}}={\mathcal{D}}(t)$ of an incompressible rigid-viscoplastic crystal for the simpified 2D model with $N=3$ slip systems. In an Eulerian description of a crystal visco-plasticity theory, the unknowns are: the velocity ${\boldsymbol{v}}:[0,T]\times{\mathcal{D}}\to\mathbb{R}^{2}$, the crystal lattice orientation, i.e. the rotation $\theta:[0,T]\times{\mathcal{D}}\to\mathbb{R}$ and the Cauchy stress ${\boldsymbol{\sigma}}:[0,T]\times{\mathcal{D}}\to\mathbb{R}^{2\times 2}_{S}$. Let ${\boldsymbol{\sigma}}={\boldsymbol{\sigma}}^{\prime}+p{\boldsymbol{I}}$, where ${\boldsymbol{\sigma}}^{\prime}$ is the stress deviator while $p:[0,T]\times{\mathcal{D}}\to\mathbb{R}$ is the pressure (mean stress).
The momentum balance in the Eulerian coordinates reads
$\rho^{mass}(\partial_{t}{\boldsymbol{v}}+{\boldsymbol{v}}\cdot{\boldsymbol{\nabla}}{\boldsymbol{v}})-\mbox{\bf div}{\boldsymbol{\sigma}}^{\prime}+\nabla p=\rho^{mass}{{\boldsymbol{f}}}\quad\hbox{ in }{\mathcal{D}},$ (2.35)
where the mass density $\rho^{mass}>0$ and the body forces ${\boldsymbol{f}}$ are supposed to be known. The incompressibility condition is
[…]
The momentum balance equations are completed by the constitutive equation, which relates the stress tensor ${\boldsymbol{\sigma}}$ and the rate of deformation tensor ${\boldsymbol{D}}({\boldsymbol{v}})$ (see (2.8)) through the evolution equations for each slip system $s$ given by (2.11) or some viscous regularization (2.13) or (2.15).
The boundary $\partial{\mathcal{D}}$ of the domain ${\mathcal{D}}$ is decomposed into two disjoint parts, $\Gamma_{v}$ and $\Gamma_{s}$, such that the velocity is prescribed on $\Gamma_{v}$ and traction is prescribed on $\Gamma_{s}$, at any time $t$:
${\boldsymbol{v}}(t)={\boldsymbol{V}}(t)\quad\hbox{on}\quad\Gamma_{v},\quad{\boldsymbol{\sigma}}(t){\boldsymbol{n}}={\boldsymbol{S}}(t)\quad\hbox{on}\quad\Gamma_{s},$ (2.37)
where ${\boldsymbol{n}}$ stands for the outward unit normal on $\partial{\mathcal{D}}$, ${\boldsymbol{V}}$ is the imposed velocity and ${\boldsymbol{S}}$ is the prescribed stress vector.
We also consider another partition of $\partial{\mathcal{D}}$ into $\partial_{in}{\mathcal{D}}(t)$ and $\partial_{out}{\mathcal{D}}(t)$ corresponding to incoming (${\boldsymbol{v}}\cdot{\boldsymbol{n}}<0$) and outcoming (${\boldsymbol{v}}\cdot{\boldsymbol{n}}\geq 0$) flux. The boundary conditions associated with the lattice evolution equations (2.19) reads
$\theta(t)=\theta^{in}(t),\quad\quad\hbox{on}\quad\partial_{in}{\mathcal{D}}(t),$ (2.38)
To the field equations, we add the initial conditions
${\boldsymbol{v}}(0)={\boldsymbol{v}}^{0},\quad\theta(0)=\theta^{0},\quad\hbox{in}\quad{\mathcal{D}}.$ (2.39)
where ${\boldsymbol{v}}^{0}$ is the initial velocity and $\theta^{0}$ gives the initial orientation of the crystal lattice. Note that in this model, there is no need to prescribe initial conditions for the stress. This is very convenient since the initial stress field is generally not known or it cannot be easily measured.
2.5 Numerical strategy
To solve the governing equations, we employ a hybrid computational approach that combines mixed finite element and discontinuous Galerkin methods, originally developed in [27]. Below, we outline the key principles underlying our numerical simulations. Our approach uses an implicit (backward) Euler time discretization scheme for the field equations, leading to a system of nonlinear equations for the velocity field ${\boldsymbol{v}}$ and the lattice orientation $\theta$. At each time step, we implement an iterative algorithm to solve these nonlinear equations using a two-part strategy: (i) the velocity field is treated using a variational formulation discretized via the finite element method, (ii) the lattice orientation is solved using a discontinuous Galerkin method with an upwind flux.
A significant advantage of this numerical framework is that it circumvents the need to model elastic deformation, as it does not rely on an elastic predictor/plastic corrector scheme, such as the one used in [36]. The rigid-viscoplastic model (2.15) poses particular computational challenges due to the non-differentiability of its viscoplastic terms, rendering standard finite element techniques for Navier–Stokes fluids [37, 38] unsuitable. We address this issue through a modified iterative decomposition–coordination formulation, coupled with the augmented Lagrangian method [39, 40]. This modification is essential because, unlike the Bingham model studied in [39, 40], the crystal model (2.11) lacks co-axiality between the stress deviator and the rate of deformation.
For simulations where the domain ${\mathcal{D}}$ evolves over time, we implement an Arbitrary Eulerian-Lagrangian (ALE) description. Throughout the numerical studies presented in subsequent sections, we employ a consistent spatial discretization scheme: P2 elements for the velocity field, P1 elements for pressure, and P1 discontinuous elements for the stress field, slip rates, dislocation densities, and lattice orientations.
During large deformations, the aspect ratio of the elements can become excessively large, which may introduce errors and hinder strain localization. To mitigate this, a multi-step mesh refinement strategy is applied at each remeshing stage, adjusting the mesh density as deformation progresses. This gradual mesh refinement enhances the simulation resolution during deformation, minimizes information loss, and avoids a significant increase in computational cost, see [41]. In our adaptive meshing approach for both monocrystal and polycrystal simulations, we apply three combined criteria: strain gradient, orientation gradient, and accumulated plastic strain.
3 Stability analysis and attractors of crystal orientation
We will suppose in this section that we deal with a velocity gradient driven processes, i.e. ${\boldsymbol{L}}={\boldsymbol{L}}^{*}$ is given and let us denote by ${\boldsymbol{D}}^{*}=({\boldsymbol{L}}^{*}+{\boldsymbol{L}}^{*T})/2$ the driven rate of deformation and by ${\boldsymbol{W}}^{*}=({\boldsymbol{L}}^{*}-{\boldsymbol{L}}^{*T})/2$ the driven rotation tensor. Bearing in mind that ${\boldsymbol{M}}_{s}={\boldsymbol{R}}{\boldsymbol{M}}_{s}^{0}{\boldsymbol{R}}^{T}$ then the slip rate decomposition of ${\boldsymbol{D}}^{*}$ (2.8) can be rewritten as ${\boldsymbol{R}}^{T}{\boldsymbol{D}}^{*}{\boldsymbol{R}}=\sum_{s=1}^{N}\dot{\gamma}^{s}{\boldsymbol{M}}_{s}^{0}$. Hence, for a given strain rate ${\boldsymbol{D}}^{*}$, the plastic (or viscoplastic) constitutive law (2.17) will give the slip rates $\dot{\gamma}^{s}$ as function of ${\boldsymbol{R}}$ only, i.e.
$(\dot{\gamma}^{1}({\boldsymbol{R}}),\dot{\gamma}^{2}({\boldsymbol{R}}),...,\dot{\gamma}^{N}({\boldsymbol{R}}))=\arg\min_{{\boldsymbol{R}}^{T}{\boldsymbol{D}}^{*}{\boldsymbol{R}}=\sum_{s=1}^{N}g^{s}{\boldsymbol{M}}_{s}^{0}}J(g^{1},g^{2},...,g^{N}).$ (3.1)
We deduce that the differential equation(2.9) is now a Cauchy problem for the rotation tensor ${\boldsymbol{R}}(t)$:
[…]
The main objective of this section is to give a framework of the stability analysis of the above differential system and to get some insights of possible attractors of the lattice rotation. The fact that the rotation group ${\mathcal{O}}=\{{\boldsymbol{Q}}\;;\;{\boldsymbol{Q}}^{T}={\boldsymbol{Q}}^{-1}\}$ is not a vectorial space is an important difficulty which occurs when we would like to use the linear stability theory to distinguish between the stable and unstable stationary rotations. That is why it could be more convenient to decompose the lattice rotation ${\boldsymbol{R}}$ in three elementary rotations using three angles ${\boldsymbol{\theta}}=(\theta_{1},\theta_{2},\theta_{3})$, and denote it by ${\boldsymbol{R}}({\boldsymbol{\theta}})$. If we deal with Tait–Bryan angles $(\theta_{1},\theta_{2},\theta_{3})=(\alpha,\beta,\gamma)$ then ${\boldsymbol{R}}({\boldsymbol{\theta}})={\boldsymbol{R}}(\theta_{1}-\theta_{1}^{0},{\boldsymbol{e}}_{1}){\boldsymbol{R}}(\theta_{2}-\theta_{2}^{0},{\boldsymbol{e}}_{2}){\boldsymbol{R}}(\theta_{3}-\theta_{3}^{0},{\boldsymbol{e}}_{3})$, where we have denoted by ${\boldsymbol{R}}(\theta_{i},{\boldsymbol{e}}_{i})$ the rotation with angle $\theta_{i}$ around the axis $Ox_{i}$ and by ${\boldsymbol{\theta}}^{0}=(\theta_{1}^{0},\theta_{2}^{0},\theta_{3}^{0})$ the initial orientation of the crystal with respect to $Ox_{1}x_{2}x_{3}$ Eulerian coordinates. If we deal with proper Euler angles $(\theta_{1},\theta_{2},\theta_{3})=(\varphi,\theta,\psi)$ then ${\boldsymbol{R}}({\boldsymbol{\theta}})={\boldsymbol{R}}(\theta_{1}-\theta_{1}^{0},{\boldsymbol{e}}_{3}){\boldsymbol{R}}(\theta_{2}-\theta_{2}^{0},{\boldsymbol{e}}_{1}){\boldsymbol{R}}(\theta_{3}-\theta_{3}^{0},{\boldsymbol{e}}_{3})$. In all cases $\dot{\gamma}^{s}$ are functions of ${\boldsymbol{\theta}}$ through
$(\dot{\gamma}^{1}({\boldsymbol{\theta}}),\dot{\gamma}^{2}({\boldsymbol{\theta}}),...,\dot{\gamma}^{N}({\boldsymbol{\theta}}))=\arg\min_{{\boldsymbol{R}}({\boldsymbol{\theta}})^{T}{\boldsymbol{D}}^{*}{\boldsymbol{R}}({\boldsymbol{\theta}})=\sum_{s=1}^{N}g^{s}{\boldsymbol{M}}_{s}^{0}}J(g^{1},g^{2},...,g^{N}).$ (3.3)
If we denote by ${\boldsymbol{\Omega}}=\dot{{\boldsymbol{R}}}{\boldsymbol{R}}^{T}$ the rotation rate tensor then from (3.2) we get:
${\boldsymbol{\Omega}}(t)={\boldsymbol{W}}^{*}-\sum_{s=1}^{N}\dot{\gamma}^{s}({\boldsymbol{\theta}}(t)){\boldsymbol{Q}}_{s}.$
Since the above equation involves only skew tensors it is more useful to use the vector representation of a skew tensor. For a skew tensor ${\boldsymbol{A}}$ let ${\boldsymbol{a}}$ be such that ${\boldsymbol{A}}{\boldsymbol{r}}={\boldsymbol{a}}\wedge{\boldsymbol{r}}$ for all ${\boldsymbol{r}}$, or $a_{1}=A_{32},a_{2}=A_{13},a_{3}=A_{21}$. If we denote by ${\boldsymbol{\omega}},{\boldsymbol{w}}^{*},{\boldsymbol{q}}_{s},{\boldsymbol{q}}_{s}^{0}$ the vectors associated to ${\boldsymbol{\Omega}},\boldsymbol{W}^{*},{\boldsymbol{Q}}_{s},{\boldsymbol{Q}}_{s}^{0}$ (note that ${\boldsymbol{q}}_{s}={\boldsymbol{q}}_{s}({\boldsymbol{\theta}})={\boldsymbol{R}}({\boldsymbol{\theta}}){\boldsymbol{q}}_{s}^{0}$) then the above equation reads
${\boldsymbol{\omega}}(t)={\boldsymbol{w}}^{*}-\sum_{s=1}^{N}\dot{\gamma}^{s}({\boldsymbol{\theta}}(t)){\boldsymbol{q}}_{s}({\boldsymbol{\theta}}(t)).$ (3.4)
There exists a basis $\{{\boldsymbol{b}}_{1},{\boldsymbol{b}}_{2},{\boldsymbol{b}}_{3}\}$, which could depend on ${\boldsymbol{\theta}}$, such that the angular velocity vector ${\boldsymbol{\omega}}$ can be written as ${\boldsymbol{\omega}}=\dot{\theta}_{1}{\boldsymbol{b}}_{1}+\dot{\theta}_{2}{\boldsymbol{b}}_{2}+\dot{\theta}_{3}{\boldsymbol{b}}_{3}$. For instance, for the Euler angles ${\boldsymbol{b}}_{1}={\boldsymbol{e}}_{3},{\boldsymbol{b}}_{2}={\boldsymbol{e}}_{N}=\cos(\varphi){\boldsymbol{e}}_{1}+\sin(\varphi){\boldsymbol{e}}_{2}$ and ${\boldsymbol{b}}_{3}=\sin(\varphi)\sin(\theta){\boldsymbol{e}}_{1}-\cos(\varphi)\sin(\theta){\boldsymbol{e}}_{2}+\cos(\theta){\boldsymbol{e}}_{3}$. Bearing in mind these notations equation (3.4) can be written now as a Cauchy problem in terms of angles ${\boldsymbol{\theta}}$ as
[…]
where the three components ${\mathcal{F}}_{i}$ of ${\boldsymbol{\mathcal{F}}}$ are given by:
$\sum_{i=1}^{3}{\mathcal{F}}_{i}({\boldsymbol{\theta}}){\boldsymbol{b}}_{i}({\boldsymbol{\theta}})={\boldsymbol{w}}^{*}-\sum_{s=1}^{N}\dot{\gamma}^{s}({\boldsymbol{\theta}}){\boldsymbol{q}}_{s}({\boldsymbol{\theta}}).$
To describe the lattice orientation attractors, we first need to compute the stationary (or invariant) lattice orientation, denoted by ${\boldsymbol{\theta}}^{st}$, as the solution to the following system:
${\boldsymbol{w}}^{*}=\sum_{s=1}^{N}\dot{\gamma}^{s}({\boldsymbol{\theta}}^{st}){\boldsymbol{q}}_{s}({\boldsymbol{\theta}}^{st}).$ (3.6)
Since the attractors are linearly stable stationary lattice orientations, the stability of ${\boldsymbol{\theta}}^{st}$ can be characterized by the eigenvalues $\lambda_{i}({\boldsymbol{\theta}}^{st})$, with $i=1,2,3$, of the $3\times 3$ matrix:
$T_{ij}({\boldsymbol{\theta}}^{st}=\frac{\partial{\mathcal{F}}_{i}}{\partial\theta_{j}}({\boldsymbol{\theta}}^{st}).$
If
$Re(\lambda_{i}({\boldsymbol{\theta}}^{st}))<0\quad\mbox{for all}\;i=1,2,3$ (3.7)
then ${\boldsymbol{\theta}}^{st}$ is linearly stable and is an attractor, denoted by ${\boldsymbol{\theta}}^{att}$. That means that there exists a neighborhood ${\mathcal{B}}$ of ${\boldsymbol{\theta}}^{att}$, called attraction basin, such that if the initial orientation ${\boldsymbol{\theta}}^{0}\in{\mathcal{B}}$ then ${\boldsymbol{\theta}}(t)\to{\boldsymbol{\theta}}^{att}$ for $t\to\infty$.
We must mention here that the application of the above stability analysis to a specific 3-D crystal (FCC, HCP, etc.) is quite complex. Indeed, it is very difficult to find the $N$ slip rate functions $\dot{\gamma}^{s}$ for the three angles ${\boldsymbol{\theta}}$ from (3.3) for a given rate of deformation ${\boldsymbol{D}}^{*}$. The second difficulty arises from solving the nonlinear system (3.6). The 3-D case is beyond the scope of this paper, but it could be addressed in future work. In the next section, we will apply the same methodology to approach the 2-D case with three slip systems, where we were able to resolve the two main difficulties outlined earlier.
4 Stability analysis of the 2-D model
Applying the above stability analysis to determine the lattice orientation attractors is a challenging task. Indeed, deriving the expression for ${\boldsymbol{\mathcal{F}}}$ and solving the nonlinear system (3.6) becomes very complex in the general case with numerous slip systems. However, for the 2-D model with three slip systems, many simplifications occur. Firstly, we deal with a single angle $\theta$ to describe the lattice orientation, and we have an analytical decomposition (2.30) instead of a minimization problem (3.1).
4.1 Lattice orientation’s Cauchy problem
Let denote by $d^{*}>0$ and $-d^{*}$ the principal rates of deformation, by ${\boldsymbol{d}}_{1}^{*},{\boldsymbol{d}}_{2}^{*}$ the principal directions (eigenvectors) of ${\boldsymbol{D}}^{*}$, by $\omega^{*}$ the Eulerian rotation rate and let $\psi^{*}=\psi^{*}(t)$ be the angle between ${\boldsymbol{d}}_{1}^{*}$ and $Ox_{1}$ axis, i.e. we have
${\boldsymbol{D}}^{*}=d^{*}({\boldsymbol{d}}_{1}^{*}\otimes{\boldsymbol{d}}_{1}^{*}-{\boldsymbol{d}}_{2}^{*}\otimes{\boldsymbol{d}}_{2}^{*}),\quad{\boldsymbol{W}}^{*}=\omega^{*}({\boldsymbol{e}}_{1}\otimes{\boldsymbol{e}}_{2}-{\boldsymbol{e}}_{2}\otimes{\boldsymbol{e}}_{1}),\quad{\boldsymbol{L}}^{*}={\boldsymbol{D}}^{*}+{\boldsymbol{W}}^{*}.$
Let us deduce here ${\mathcal{F}}(\theta)$ the expressions of ${\boldsymbol{\mathcal{F}}}$ for the 2D model. For that let us remark that from (2.33) we get
$\sum_{s=1}^{3}\dot{\gamma}_{s}=d^{*}{\mathcal{G}}(\theta-\psi^{*})$ (4.1)
where ${\mathcal{G}}$ is given in (2.34) and plotted in Figure 1e. Finally, from (2.19) we deduce ${\mathcal{F}}(\theta)=\frac{1}{2}(d^{*}{\mathcal{G}}(\theta-\psi^{*})-2\omega^{*})$ and the 2D expression of (3.5) :
[…]
In the following, we consider two types of Eulerian strain-driven processes. The first involves a constant, i.e., stationary, velocity gradient with $\psi^{(}t)=\text{const}$. The second, referred to here as the "Lagrangian stationary velocity gradient," has $\psi^{(}t)=\text{const}-\omega^{*}t$. In this case, the directions of the Eulerian strain gradient relative to the lattice remain constant over time.
4.2 Stability analysis for stationary velocity gradients
We will suppose here that the angle $\psi^{*}(t)$ of the principal directions of ${\boldsymbol{D}}^{*}$ is constant in time, i.e.
$\psi^{*}(t)=\psi^{*}_{0},$
and we denote by $\tilde{\theta}(t)$ the crystal orientation with respect to ${\boldsymbol{d}}^{*}_{1}$ axis:
$\tilde{\theta}(t)=\theta(t)-\psi^{*}_{0}.$
Bearing in mind this notation (4.2) reads
[…]
We can compute now the stationary orientations $\tilde{\theta}^{st}$ as the solution of the nonlinear algebraic equation
${\mathcal{G}}(\tilde{\theta}^{st})=2\frac{\omega^{*}}{d^{*}}.$ (4.4)
Let $\tilde{\theta}^{st}$ be a stationary orientation and suppose that ${\mathcal{G}}$ is differentiable in $\tilde{\theta}^{st}$. Then
$\mbox{if}\quad{\mathcal{G}}^{\prime}(\tilde{\theta}^{st})<0\quad\mbox{then }\;\tilde{\theta}^{st}\;\mbox{is an attractor.}$. (4.5)
In Fig. 2 we have plotted the function $\theta\to{\mathcal{G}}(\theta)$ in the case of in-plane deformation of a FCC crystal. We remark that this function is piece-wise linear and, as it follows from (4.4), the stationary lattice orientation $\tilde{\theta}^{st}$ have to be founded at the intersection of this graph with the horizontal line $y=2\omega/d$.
Figure 2: The function $\theta\mapsto\mathcal{G}(\theta)$ in the case of in-plane deformation of an FCC crystal. Its intersection with the line $y=2\omega/d$ defines the stationary lattice orientations $\tilde{\theta}^{\text{st}}$. The attractors $\tilde{\theta}^{\text{att}}$ are shown as solid colored discs (orange, blue, and green), and the colored segments (matching the color of each attractor) represent their respective basins of attraction.
In what follows we have chosen the time scale to be $1/d^{*}$ such that the non-dimensional reference strain $\gamma=d^{*}t$ has the same values as the time $t$. We can distinguish 3 cases :
i) $|\omega^{*}|<d^{*}$. In this case, plotted in Figure 2a, there are 6 stationary orientations $\tilde{\theta}^{st})$ given by
$\tilde{\theta}_{1}^{st}=\frac{\omega^{*}}{d^{*}}\frac{\phi}{4},\quad\tilde{\theta}_{2}^{st}=\frac{\phi}{4}+(1-\frac{\omega^{*}}{d^{*}})\frac{\pi-\phi}{8},\quad\tilde{\theta}_{3}^{st}=\frac{\pi}{4}+(1+\frac{\omega^{*}}{d^{*}})\frac{\pi-\phi}{8},$
$\tilde{\theta}_{4}^{st}=\frac{\pi}{2}-\frac{\phi}{4}+(1-\frac{\omega^{*}}{d^{*}})\frac{\phi}{4},\quad\tilde{\theta}_{5}^{st}=\frac{\pi}{2}+\frac{\phi}{4}+(1+\frac{\omega^{*}}{d^{*}})\frac{\pi-\phi}{8},\quad\tilde{\theta}_{6}^{st}=\frac{3\pi}{4}+(1-\frac{\omega^{*}}{d^{*}})\frac{\pi-\phi}{8}$
but following (4.5) only three ($\tilde{\theta}_{2}^{st},\tilde{\theta}_{4}^{st}$ and $\tilde{\theta}_{6}^{st}$) are attractors (stable):
$\tilde{\theta}_{1}^{att}=\frac{\phi}{4}+(1-\frac{\omega^{*}}{d^{*}})\frac{\pi-\phi}{8},\;\tilde{\theta}_{2}^{att}=\frac{\pi}{2}-\frac{\phi}{4}+(1-\frac{\omega^{*}}{d^{*}})\frac{\phi}{4},\;\tilde{\theta}_{3}^{att}=\frac{3\pi}{4}+(1-\frac{\omega^{*}}{d^{*}})\frac{\pi-\phi}{8},$ (4.6)
which are plotted in Figure 2 by solid colored (orange, blue and green) discs. Their attraction basins are the intervals plotted in Fig. 2 by colored segments (corresponding to the color of the attractor) given by
$B^{att}_{1}=(\tilde{\theta}_{1}^{st},\tilde{\theta}_{3}^{st}),\quad B^{att}_{2}=(\tilde{\theta}_{3}^{st},\tilde{\theta}_{5}^{st}),\quad B^{att}_{3}=(\tilde{\theta}_{5}^{st},\pi+\tilde{\theta}_{1}^{st}).$
This means that
$\mbox{if}\quad\tilde{\theta}_{0}\in B^{att}_{i}\quad\mbox{then }\quad\tilde{\theta}(t)\to\tilde{\theta}_{i}^{att}\quad\mbox{for }\quad t\to\infty,\quad\mbox{for all}\;i=1,2,3.$ (4.7)
In Fig. 3a, we plot the numerical simulation of the time evolution of the lattice orientations, $t\to(\cos(\theta(t)),\sin(\theta(t)))$, for 200 different initial orientations $\theta_{0}\in[0,2\pi]$ and $\psi^{*}_{0}=0$. We have chosen $\omega^{=}d^{*}/2$, corresponding to case 1 of Fig. 2. We observe the presence of 3 different attractors (6 in Fig. 3a, with $\pi$ symmetry), as described by the theoretical results above. Note that the attractors are only valid for large strains, i.e., the distance $|\theta(t)-\theta_{i}^{att}|$ is small enough only for strains larger than $50\%$. These large strains occur frequently, even for small overall strains.
Figure 3: Numerical simulation of 200 time/strain trajectories $t\to(t,\cos(\theta(t)),\sin(\theta(t)))$ of the crystal orientation for different choices of the initial orientation $\theta_{0}$. (a): case 1 ($|\omega^{*}|<d^{*}$ and $\psi^{*}_{0}=0$) with 3 attractors. (b): case 2 ($-\omega^{*}=d^{*}$ and $\psi_{0}^{*}=0$) with 3 half-attractors. (c): case 3 ($|\omega|>d$ and $\psi^{*}_{0}=0$) with no attractors.
ii) $|\omega^{*}|=d^{*}$. In this case, plotted in Fig. 2, which is specific to the slip bands (see section 5.3), there are 3 stationary orientations $\tilde{\theta}$ given by
$\tilde{\theta}_{1}^{st}=\mbox{sign}(\frac{\omega^{*}}{d^{*}})\frac{\phi}{4},\quad\tilde{\theta}_{2}^{st}=\frac{\pi}{2}+\mbox{sign}(\frac{\omega^{*}}{d^{*}})\frac{\pi}{4},\quad\tilde{\theta}_{3}^{st}=\frac{\pi}{2}-\mbox{sign}(\frac{\omega^{*}}{d^{*}})\frac{\phi}{4}.$
For all these angles the function $\theta\to\ {\mathcal{G}}(\theta)$ is not differentiable but the left and right derivative exists and have opposite signs. Following (4.5) neither are stable (in the classical sense) but they have left or right attractors, that is why be will call them "half-attractors". Indeed, for $\omega^{*}=d^{*}$ we have
$\tilde{\theta}_{3}^{att}=\frac{\phi}{4},\quad\tilde{\theta}_{2}^{att}=\frac{\pi}{2}-\frac{\phi}{4},\quad\tilde{\theta}_{1}^{att}=\frac{3\pi}{4},\quad\mbox{for}\;\omega^{*}=d^{*},$ (4.8)
$B^{att}_{3}=[\frac{\phi}{4},\frac{\pi}{2}-\frac{\phi}{4}),\quad B^{att}_{2}=[\frac{\pi}{2}-\frac{\phi}{4},\frac{3\pi}{4}),\quad B^{att}_{1}=[\frac{3\pi}{4},\pi+\frac{\phi}{4}),\quad\mbox{for}\;\omega^{*}=d^{*}.$
For $d^{*}=-\omega^{*}$ the attractors
$\tilde{\theta}_{1}^{att}=\frac{\pi}{4},\quad\tilde{\theta}_{3}^{att}=\frac{\pi}{2}+\frac{\phi}{4},\quad\tilde{\theta}_{2}^{att}=\pi-\frac{\phi}{4},\quad\mbox{for}\;\omega^{*}=-d^{*},$ (4.9)
are plotted in Fig. 2 by solid colored (orange, blue and green) discs, while their attraction basins
$B^{att}_{1}=(-\frac{\phi}{4},\frac{\pi}{4}],\quad B^{att}_{3}=(\frac{\pi}{4},\frac{\pi}{2}+\frac{\phi}{4}],\quad B^{att}_{2}=(\frac{\pi}{2}+\frac{\phi}{4},\pi-\frac{\phi}{4}],\quad\mbox{for}\;\omega^{*}=-d^{*},$
are the intervals plotted by colored segments (corresponding to the color of the attractor).
For $d^{*}=-\omega^{*}$ (corresponding to case 2 of Fig. 2) we have computed the time evolution of the lattice orientations $t\to(\cos(\theta(t)),\sin(\theta(t)))$ for 200 choices of the initial orientation $\theta_{0}\in[0,2\pi)$ (see Fig. 3b). We remark the presence of 3 attractors and the lattice orientations trajectories obey (4.7). As before the attractors work only for large strains.
We have to mention here that in the case $|\omega^{*}|=d^{*}$ the attractors orientation correspond to the specific orientations where only one system is active (see section 2.3.4). The label $r$ of each attractor $\tilde{\theta}_{r}^{att}$ was chosen such that only the system $r$ is active (i.e. $\dot{\gamma}^{r}\neq 0$ and $\dot{\gamma}^{s}=0$ for $s\neq r$).
iii) $|\omega|>d$. In this case, plotted in Fig. 2, there are no stationary orientations. To see that we have chosen $\omega^{*}=\frac{3}{2}d^{*}=1.5$ (corresponding to case 3 of Fig. 2) and we have plotted in Fig. 3c the numerical simulation of the time evolution of the lattice orientations $t\to(\cos(\theta(t)),\sin(\theta(t)))$ for 200 choices of the initial orientation $\theta_{0}\in[0,2\pi]$ and $\psi^{*}_{0}=0$. We remark the absence of any attractor.
4.3 Stability analysis for Lagrangian stationary velocity gradients
Let us suppose here that the principal directions of the Eulerain strain rate ${\boldsymbol{D}}^{*}$ are rotated with the angle $-\omega^{*}t$, such that it has the same orientation after the rigid rotation imposed by the spin tensor ${\boldsymbol{W}}^{*}$. That is why we will suppose that the angle $\psi^{*}(t)$ is given by
$\psi^{*}(t)=\psi^{*}_{0}-\omega^{*}t,$
and we define the "rotated" crystal orientation to be
$\tilde{\theta}(t)=\theta(t)-\psi^{*}_{0}+\omega^{*}t.$
This situation arrives for the rotation and a radial expansion deformation of the type ${\boldsymbol{v}}^{*}=\frac{V^{*}R^{*}}{r}{\boldsymbol{e}}_{r}-\omega^{*}r{\boldsymbol{e}}_{\varphi}$ where $r,\varphi$ are the Eulerian cylindrical coordinates. In this case ${\boldsymbol{d}}_{1}={\boldsymbol{e}}_{\varphi},{\boldsymbol{d}}_{2}={\boldsymbol{e}}_{r}$ and $\psi^{*}(t)=\pi/2+\varphi-\omega^{*}t$.
Bearing in mind this notation (4.2) reads
[…]
which is exactly (4.3) with $\omega^{*}=0$. We can deduce that there are 6 stationary orientations $\tilde{\theta}^{st}$ given by
$\tilde{\theta}_{1}^{st}=0,\quad\tilde{\theta}_{2}^{st}=\frac{\pi+\phi}{8},\quad\tilde{\theta}_{3}^{st}=\frac{3\pi-\phi}{8},\quad\tilde{\theta}_{4}^{st}=\frac{\pi}{2},\quad\tilde{\theta}_{5}^{st}=\frac{3\pi+\phi}{8},\quad\tilde{\theta}_{6}^{st}=\frac{7\pi-\phi}{8}$
but following (4.5) only three ($\tilde{\theta}_{2}^{st},\tilde{\theta}_{4}^{st}$ and $\tilde{\theta}_{6}^{st}$) are attractors (stable):
$\tilde{\theta}_{1}^{att}=\frac{\pi+\phi}{8},\;\tilde{\theta}_{2}^{att}=\frac{\pi}{2},\;\tilde{\theta}_{3}^{att}=\frac{7\pi-\phi}{8},$ (4.11)
Their attraction basins are the intervals given by
$B^{att}_{1}=(\tilde{\theta}_{1}^{st},\tilde{\theta}_{3}^{st}),\quad B^{att}_{2}=(\tilde{\theta}_{3}^{st},\tilde{\theta}_{5}^{st}),\quad B^{att}_{3}=(\tilde{\theta}_{5}^{st},\pi+\tilde{\theta}_{1}^{st}),$
which means that
$\mbox{if}\quad\tilde{\theta}_{0}\in B^{att}_{i}\quad\mbox{then }\quad\tilde{\theta}(t)\to\tilde{\theta}_{i}^{att}\quad\mbox{for }\quad t\to\infty,\quad\mbox{for all}\;i=1,2,3.$ (4.12)
5 Numerical simulations
In this section we present some numerical simulations of the crystal lattice orientation evolution for three different classical problems in crystal plasticity. The first one deals with a poly-crystal (i.e. a non-homogeneous initial orientation) which is loaded with an homogeneous gradient velocity loading. The second one concerns a mono-crystal which is subjected to an isotropic loading, as a void growth under radial expansion. In this case the associated rate of deformation ${\boldsymbol{D}}^{*}$ is radial, hence non-homogeneous in space while the initial orientation is homogeneous. The third one deals with a slip band deformation in a uni-axial loading of a mono-crystal.
We have to mention here that in all cases only the boundary conditions are given in relation to a driven velocity field ${\boldsymbol{v}}^{*}$ but the computed velocity field ${\boldsymbol{v}}$ is different from the expected one. Indeed the response of the (poly-)crystal on the boundary loading, could presents slip and/or kink bands, which cannot be described by the driven velocity field ${\boldsymbol{v}}^{*}$. However, in the examples shown here, the overall final orientation of the lattice can be deduced from the apriori attractors distribution predicted by the stability analysis of the previous section.
The material parameter used in this quasi-static computation is not specific to a particular material but instead reflects the order of magnitude typically associated with hexagonal close-packed (HCP) crystals ($\phi=\pi/3=60^{\circ}$) for the first two simulations and with a face-centered cubic (FCC) crystal ($\phi=54.7^{\circ}$) for the third. The material parameters are as follows: density $\rho^{\text{mass}}=3200,\text{kg/m}^{3}$, yield limit $\tau_{c}^{s}=20,\text{MPa}$ for HCP, and $\rho^{\text{mass}}=16650,\text{kg/m}^{3}$, $\tau_{c}^{s}=220,\text{MPa}$ for FCC. In all cases, no hardening effects were considered, and a small viscosity (less than $1\%$ of $\eta^{\text{ref}}=\tau_{c}^{s}/\dot{\Gamma}_{c}$, where $\dot{\Gamma}_{c}$ is the characteristic strain rate) was used for numerical reasons.
5.1 Poly-crystal Response to Homogenenous Loading
The poly-crystal domain is a rectangle $(0,L(t))\times((0,H(t))$, which is a square ($L_{0}=H_{0}$) at the initial configuration $t=0$, containing 15 grains of different orientations (see Fig. 4 right).
Figure 4: Left: Schematic representation of a polycrystal subjected to a homogeneous velocity gradient loading. Right: Spatial distribution $(x_{1},x_{2})\mapsto\theta^{0}(x_{1},x_{2})$ of the initial lattice orientation, shown in a color scale ranging from $-30^{\circ}$ to $30^{\circ}$.
Figure 5: Top: Computed Eulerian distribution $(x_{1},x_{2})\mapsto\theta(t,x_{1},x_{2})$ of the lattice orientation, shown in a color scale ranging from $-30^{\circ}$ to $30^{\circ}$. Middle: Expected Eulerian distribution $(x_{1},x_{2})\mapsto\theta^{\text{att}}(t,x_{1},x_{2})$ of the orientation attractor, in the same color scale $(-30^{\circ},30^{\circ})$. Bottom: Eulerian distribution $(x_{1},x_{2})\mapsto|\theta(t,x_{1},x_{2})-\theta^{\text{att}}(x_{1},x_{2})|$ representing the difference between the lattice orientation and the estimated attractor, shown in a color scale from $0^{\circ}$ to $30^{\circ}$. The distributions are shown for four time points: $t=0$ ($\epsilon^{\text{eng}}=0$), $t=T/3$ ($\epsilon^{\text{eng}}=0.166$), $t=2T/3$ ($\epsilon^{\text{eng}}=0.33$), and $t=T$ ($\epsilon^{\text{eng}}=0.5$).
On the boundary, denoted by $\Gamma_{v}(t)=\Gamma(t)$, we impose a velocity ${\boldsymbol{V}}$ (see Fig. 4 left), corresponding to an in-compressible Eulerian velocity field ${\boldsymbol{v}}^{*}=d^{*}(x_{1}{\boldsymbol{e}}_{1}-x_{2}{\boldsymbol{e}}_{2})$, i.e.
${{\boldsymbol{V}}}(t)(x_{1},x_{2})=d^{*}(x_{1}{\boldsymbol{e}}_{1}-x_{2}{\boldsymbol{e}}_{2}).$ (5.1)
The associated gradient of the velocity field ${\boldsymbol{v}}^{*}$ is
${\boldsymbol{L}}^{*}={\boldsymbol{D}}^{*}=d^{*}({\boldsymbol{e}}_{1}\otimes{\boldsymbol{e}}_{1}-{\boldsymbol{e}}_{2}\otimes{\boldsymbol{e}}_{2}),$
which is homogeneous and corresponds to the choice
${\boldsymbol{d}}_{1}^{*}={\boldsymbol{e}}_{1},\quad{\boldsymbol{d}}_{2}^{*}={\boldsymbol{e}}_{2},\quad\omega^{*}=0,\quad\psi^{*}=\psi_{0}^{*}=0.$
Following the stability analysis of the rotated orientation angle $\tilde{\theta}=\theta$ we deduce that the orientation attractor (plotted in Fig. 5 middle) is given on each grain depending on the initial orientation $\theta^{0}$ (plotted in Fig. 4 right) by
$\theta^{att}(x,y)=\left\{\begin{aligned} \frac{\pi}{6}&\quad\mbox{if}\;&\theta^{0}(x,y)&\in(0,\frac{\pi}{3})\\ -\frac{\pi}{6}&\quad\mbox{if}\;&\theta^{0}(x,y)&\in(-\frac{\pi}{3},0).\end{aligned}\right.$
The initial velocity and the initial crystal orientation where chosen to be
${\boldsymbol{v}}^{0}(x_{1},x_{2})={\boldsymbol{v}}^{*}(x_{1},x_{2}),\quad\theta^{0}=\theta^{0}(x_{1},x_{2}).$ (5.2)
Denoting by $\epsilon^{eng}$ the engineering deformation
$\epsilon^{eng}(t)=\frac{H_{0}-H(t)}{H_{0}},$
the time period $[0,T]$ and the strain rate $d^{*}=\dot{\Gamma}_{c}$ were chosen to correspond to a final engineering deformation $\epsilon^{eng}_{final}=(H_{0}-H_{final})/H_{0}=0.5=50\%$ (here $H(t)$ is the actual height of the poly-crystal).
Figure 6: Evolution of the $L^{2}$ norm $||g(t)||_{2}$ of the gap between the lattice orientation and the estimated attractor as a function engineering deformation $\epsilon^{eng}$.
In Fig. 5 (top) we have plotted the evolution of lattice orientation $(x_{1},x_{2})\to\theta(t,x_{1},x_{2})$ during the deformation process. We observe that globally the lattice orientation approaches the estimated attractor $(x_{1},x_{2})\to\theta^{\text{att}}(x_{1},x_{2})$ (plotted in Fig. 5 middle) for large strains. The final orientation does not match perfectly with the predicted one but remains very close (less than $5^{\circ}$ almost everywhere). This can be quantified by examining the gap between the lattice orientation and the estimated attractor:
$(x_{1},x_{2})\to g(t,x_{1},x_{2})=|\theta(t,x_{1},x_{2})-\theta^{\text{att}}(x_{1},x_{2})|,$
plotted in Fig. 5 (bottom). To establish a global quantitative measure, we have computed the $L^{2}$ norm of the gap $g(t)$ given by:
[…]
and plotted it in Fig. 6. We observe that the $L^{2}$ norm shows a significant decrease until $\epsilon^{\text{eng}}=0.3$, after which it exhibits a plateau around $0.06~{}\text{rad}\approx 3.5^{\circ}$. The small deviations in certain grains can be explained by the loss of strain rate homogeneity and the presence of shear bands, which initially follow grain boundaries but later appear to be associated with the overall geometry of the sample.
The initial choice of grain orientations, shown in Fig. 4 (right), includes only grain orientations that belong to an attraction basin. To explore what happens if a grain has an unstable stationary orientation, we changed the orientation of the upper-right grain (see Fig. 7) to $\theta=0$, which is an unstable stationary orientation located between two attraction basins, $(-\pi/3,0)$ and $(0,\pi/3)$. We observe that part of the grain joins the $\theta^{att}=\pi/6$ attractor, another part joins the $\theta^{att}=-\pi/6$ attractor, while a third part of the grain exhibits orientation bands associated with kink or shear bands. It is important to note that computations in this unstable regime may be mesh or time-step dependent, making it difficult to predict the final configuration.
Figure 7: Computed Eulerian distribution $(x,y)\mapsto\theta(t,x,y)$ of the lattice orientation, shown in a color scale ranging from $-30^{\circ}$ to $30^{\circ}$, for the following times: $t=0$ ($\epsilon^{\text{eng}}=0$), $t=T/3$ ($\epsilon^{\text{eng}}=0.166$), $t=2T/3$ ($\epsilon^{\text{eng}}=0.33$), and $t=T$ ($\epsilon^{\text{eng}}=0.5$). Note that one grain has an initial unstable orientation.
5.2 Mono-crystal Response to Heterogeneous Loading
At the initial state, the crystal domain is represented by a disc of radius $R_{0}$, containing an initially circular void at its center with radius $r_{0}=R_{0}/10$ (see Fig. 8, left). On the external boundary, denoted by $\Gamma_{e}(t)$, we impose a radial velocity ${\boldsymbol{V}}$ corresponding to an incompressible Eulerian velocity field,
${{\boldsymbol{V}}}(t)(x_{1},x_{2})=\frac{V_{0}R_{0}}{x_{1}^{2}+x_{2}^{2}}(x_{1}{\boldsymbol{e}}_{1}+x_{2}{\boldsymbol{e}}_{2}),$ (5.3)
where $r$ and $\varphi$ are the cylindrical coordinates. The internal boundary, denoted by $\Gamma_{i}(t)$, is traction-free (i.e., ${{\boldsymbol{S}}}(t)=0$).
Figure 8: Left: Schematic representation of a single crystal with a circular void at its center, subjected to radial velocity loading. Right: Distribution $(x,y)\mapsto\theta^{\text{att}}(x,y)$ of the predicted attractor on the initial configuration, shown in a radian color scale ranging from $-\pi/6$ to $\pi/6$.
The associated gradient of the velocity field ${\boldsymbol{v}}^{*}$ is
${\boldsymbol{L}}^{*}={\boldsymbol{D}}^{*}=d^{*}(r)({\boldsymbol{e}}_{\varphi}\otimes{\boldsymbol{e}}_{\varphi}-{\boldsymbol{e}}_{r}\otimes{\boldsymbol{e}}_{r}),\quad d^{*}(r)=\frac{V_{0}R_{0}}{r^{2}},$
which is non-homogeneous and corresponds to the choice
${\boldsymbol{d}}_{1}^{*}={\boldsymbol{e}}_{\varphi}(\varphi),\quad{\boldsymbol{d}}_{2}^{*}={\boldsymbol{e}}_{r}(\varphi),\quad\omega^{*}=0,\quad\psi^{*}=\psi_{0}^{*}=\varphi+\frac{\pi}{2}.$ (5.4)
Following the stability analysis of the rotated orientation angle $\tilde{\theta}=\theta-\psi_{0}^{*}$ we deduce that the orientation attractor $\theta^{att}=\tilde{\theta}^{att}+\psi_{0}^{*}$ is given on each of 6 circular sectors of the plane (see Fig. 8 right) by
$\theta^{att}(\varphi)=\varphi-\frac{(k+1)\pi}{3},\quad\mbox{for}\;\varphi\in\left(\frac{(1+2k)\pi}{6},\frac{(3+2k)\pi}{6}\right),\;k\geq 0.$ (5.5)
The initial conditions for the velocity and crystal orientation were chosen as
${\boldsymbol{v}}^{0}(x_{1},x_{2})={\boldsymbol{v}}^{*}(x_{1},x_{2}),\quad\theta^{0}=0.$ (5.6)
The time period $[0,T]$, the initial external radius $R_{0}$, and the velocity $V_{0}$ were selected to yield a final engineering volumetric strain of
$\epsilon^{eng}_{\text{final}}=\frac{2V_{0}T}{R_{0}}=0.1875=18.75\%.$ (5.7)
Note that the expected deformation at the void level is approximately $R_{0}/r_{0}=10$ times higher (around $180\%$) than at the external boundary.
Figure 9: Top: Computed Eulerian distribution $(x,y)\mapsto\theta(t,x,y)$ of the lattice orientation, shown in a radian color scale ranging from $-\pi/6$ to $\pi/6$. Bottom: Eulerian distribution $(x,y)\mapsto|\theta(t,x,y)-\theta^{\text{att}}(t,x,y)|$ of the gap between the lattice orientation and the estimated attractor, shown in a radian color scale ranging from 00 to $\pi/6$. The distributions correspond to the following times: $t=0$ ($\epsilon^{\text{eng}}=0$), $t=T/3$ ($\epsilon^{\text{eng}}=0.0625$), $t=2T/3$ ($\epsilon^{\text{eng}}=0.125$), and $t=T$ ($\epsilon^{\text{eng}}=0.1875$).
In Fig. 9 (top), we plot the evolution of the lattice orientation $(x,y)\mapsto\theta(t,x,y)$ during the deformation process. We observe that the orientation exhibits discontinuities at the angles $\varphi=\pi/6+k\pi/3$, consistent with the attractor structure. However, these discontinuities are more pronounced near the void, where deformations are significantly larger, than near the external boundary. This observation aligns with our stability analysis from the previous section: the distance between the lattice orientation and its attractor decreases substantially for large strains (greater than $50\%$), which occur near the void, but remains small at the external boundary.
In Fig. 9 (bottom), we show the distribution $(x,y)\mapsto|\theta(t,x,y)-\theta^{att}(t,x,y)|$, representing the difference between the lattice orientation and the estimated attractor. We observe that, overall, the blue regions (where the orientation is close to the attractor) expand during the deformation. However, as discussed above, there are regions—namely along the discontinuity lines $\varphi=\pi/6+k\pi/3$ and areas far from the void (shown in red)—where the attractor fails to accurately describe the lattice orientation.
We should also mention that very close to the void, at large strains, the deformation process becomes much more complex: slip and kink bands appear. Moreover, during the deformation process, the computed equivalent strain rate is no longer radial, as expected from ${\boldsymbol{D}}^{*}$. This is why the previous analysis, which relies on a simple assumption about the velocity field ${\boldsymbol{v}}^{*}$, is less pertinent. This can be observed in Fig. 9 (bottom), where at $t=2T/3$ ($\epsilon^{eng}=0.125$) and at $t=T$ ($\epsilon^{eng}=0.1875$), the difference between the lattice orientation and the estimated attractor is larger very close to the void.
Figure 10: Magnified view of the void region highlighted in Fig. 9. Top: Computed Eulerian distribution of lattice orientation $(x,y)\to\theta(t,x,y)$ shown in radian color scale $(-\pi/6,\pi/6)$. Bottom: Eulerian distribution of deviation between lattice orientation and estimated attractor $(x,y)\to|\theta(t,x,y)-\theta^{att}(t,x,y)|$ displayed in radian color scale $(0,\pi/6)$. Distributions are presented at four deformation stages: $t=0$ ($\epsilon^{eng}=0$), $t=T/3$ ($\epsilon^{eng}=0.0625$), $t=2T/3$ ($\epsilon^{eng}=0.125$) and $t=T$ ($\epsilon^{eng}=0.1875$).
5.3 Slip band deformation in a mono-crystal
The mono-crystal domain, which has a initial rectangular shape $(0,L)\times(0,H)$, is pulled on the upper part with a normal velocity $V$ and it is fixed on bottom. All other boundaries are stress free. We analyze one slip band localized deformation of width $H_{b}$ and orientated at the angle $\alpha$ with respect to the $Ox_{1}$ axis. In the slip band we expect the velocity distribution ${\boldsymbol{v}}^{*}(y_{1},y_{2})=V_{b}\frac{y_{2}}{H_{b}}{\boldsymbol{c}}_{1}$, where we have denoted by ${\boldsymbol{c}}_{1},{\boldsymbol{c}}_{2}$ the basis associated to the shear band and by $(y_{1},y_{2})$ the shear band variables (see Fig. 11 ), and $V_{b}=V/\sin(\alpha)$. Denoting by $2d^{*}=V_{b}/H_{b}$ we deduce that the expected velocity gradient in the shear band is
${\boldsymbol{L}}^{*}=2d^{*}{\boldsymbol{c}}_{1}\otimes{\boldsymbol{c}}_{2},\quad{\boldsymbol{D}}^{*}=d^{*}({\boldsymbol{c}}_{1}\otimes{\boldsymbol{c}}_{2}+{\boldsymbol{c}}_{2}\otimes{\boldsymbol{c}}_{1}),\quad\boldsymbol{W}^{*}=d^{*}({\boldsymbol{c}}_{1}\otimes{\boldsymbol{c}}_{2}-{\boldsymbol{c}}_{2}\otimes{\boldsymbol{c}}_{1}).$
We can compute the principal strain rate directions ${\boldsymbol{d}}_{1}^{*}=({\boldsymbol{c}}_{1}+{\boldsymbol{c}}_{2})/\sqrt{2},{\boldsymbol{d}}_{2}^{*}=({\boldsymbol{c}}_{2}-{\boldsymbol{c}}_{1})/\sqrt{2}$ to deduce that we deal with a stationary velocity gradient with $\psi^{*}=\alpha+\pi/4$.
Figure 11: Schematic description of a slip band under tensile loading.
Let use here the stability analysis developed in section 4.2 to predict the lattice orientation in a slip band in traction with $\alpha<\pi/2$. First, we remark that we deal with case ii) with $\omega^{*}=d^{*}$ where there are three half-attractors (i.e. attractors on the left but not on the right):
$\theta_{3}^{att}=\frac{\pi}{4}+\frac{\phi}{4}+\alpha,\quad\theta_{2}^{att}=\frac{3\pi}{4}-\frac{\phi}{4}+\alpha,\quad\theta_{1}^{att}=\pi+\alpha,\quad$
with 3 basins of attraction:
$B^{att}_{3}=[\frac{\pi}{4}+\frac{\phi}{4}+\alpha,\frac{3\pi}{4}-\frac{\phi}{4}+\alpha),\quad B^{att}_{2}=[\frac{3\pi}{4}-\frac{\phi}{4}+\alpha,\pi+\alpha),\quad B^{att}_{1}=[\pi+\alpha,\frac{5\pi}{4}-\frac{\phi}{4}+\alpha).$
Note that each attractor corresponds to a single active system and the basins label $s$ have been chosen to correspond to the active slip system $s$ (see Fig. 12). If we denote by $\beta_{s}$ the miss-orientation between the active slip system of the attractor $s$ and the slip band direction we get that
$\beta_{1}=0,\quad\beta_{2}=\beta_{3}=\frac{\pi}{4}-\frac{3\phi}{4}.$
We deduce that the first attractor has the active slip system parallel ($\beta_{1}=0$) with the slip band. We arrive at the same conclusion for the other two attractors ($\beta_{2}=\beta_{3}=0$) in the case of a hexagonal crystal ($\phi=\pi/3$). For a FCC crystal ($\phi=54.7^{\circ}$) there exits a slight miss-orientation $\beta_{2}=\beta_{3}=4^{\circ}$, and the slip band could be re-considered as a kink band.
In conclusion, for a slip band under traction, the stability analysis predicts that the lattice orientation of the crystal within the band will rotate in a clockwise direction, such that one of the slip systems will become parallel (or nearly so, in the case of an FCC crystal) to the slip band. It is important to note that in the above analysis, we assumed that the direction of the slip band is given and independent of the lattice rotation. However, this is not generally the case, as a more complex process is expected, where the sample shape, boundary conditions, and crystal orientation interact to explain the existence of one or more slip bands and their orientations.
Figure 12: Stability analysis for a slip band under tensile loading ($\alpha>0$): Schematic representation of three half-attractors $\theta_{s}^{att}$ with their corresponding attraction basins $B^{att}_{s}$ (green for $s=1$, blue for $s=2$, and orange for $s=3$). For each half-attractor, the single active slip direction is highlighted in bold. The parameter $\beta$ represents the misorientation between the active slip system and the slip band orientation.
For the numerical simulation, we considered an FCC crystal with an initial square shape ($L=H$), and the following initial conditions:
${\boldsymbol{v}}^{0}(x_{1},x_{2})=V\frac{x_{2}}{H}{\boldsymbol{e}}_{2},\quad\theta^{0}=0.$ (5.8)
The time period $[0,T]$ and the velocity $V$ were chosen to correspond to a final engineering deformation $\epsilon^{eng}_{final}=VT/H=0.2=20\%$. Note that the expected deformation in the slip band is expected to be much larger.
Figure 13: The computed accumulated plastic strain $\epsilon^{p}(t)$ (in color scale) and the velocity distribution (arrows) at three deformation stages: $t=T/3$ ($\epsilon^{eng}=0.067$) (left); $t=2T/3$ ($\epsilon^{eng}=0.133$) (center); and $t=T$ ($\epsilon^{eng}=0.2$) (right).
In Fig. 13 we have plotted the evolution of the deformation. We remark the presence of a slip band orientated at $\alpha=\pi/4$ where the plastic strain $\epsilon^{p}$ (i.e. $\dot{\epsilon}^{p}=|{\boldsymbol{D}}({\boldsymbol{v}})|$ ) is accumulated during the process. We remark that the deformation assumption with two rigid blocks separated by a slip band (see Fig. 11) is globally verified. Moreover, the level of plastic strain in the slip band is important (more than $2=200\%$) hence the above stability analysis could be pertinent.
As it follows from the above stability analysis (see also Fig. 12) the initial orientation of the crystal $\theta_{0}=0$ is in attraction basins $B_{2}^{att}$ with the attractor $\theta_{2}^{att}=-\phi/4$ hence the expected lattice orientation in the slip band is $-\phi/4\approx-13.7^{\circ}=0.24$rad. In Fig. 14 we have plotted the evolution of the distribution of the lattice orientation $\theta(t)$ (in rad, up) and of the its gap $\min_{s}|\theta(t)-\theta_{s}^{att}|$ with the slip band attractor (in rad, bottom). We remark that, as expected, the orientation in the slip band changes during the slip to reach asymptotically the attractor.
Figure 14: Up: the lattice orientation $\theta(t)$ (in rad ranging between $-\phi/4$ and $\phi/4$). Bottom: the gap $\min_{s}|\theta(t)-\theta_{s}^{att}|$ between the lattice orientation and the slip band attractor (in rad ranging between 00 and $\phi/4$). For $t=0$ ($\epsilon^{eng}=0$), $t=T/3$ ($\epsilon^{eng}=0.067$), $t=2T/3$ ($\epsilon^{eng}=0.133$) and $t=2T/3$ ($\epsilon^{eng}=0.2$).
Concerning the active slip system in the slip band we expect that only system $s=2$ is active in the slip band. To check this we have plotted in Fig. 15 the computed distribution of the accumulated plastic slip $\gamma^{s}_{a}$ (i.e. $\dot{\gamma}_{a}^{s}=|\dot{\gamma}^{s}|$) at the end of the process $t=T$ ($\epsilon^{eng}=0.2$). We remark that in the shear band only the system $s=2$ is active.
6 Conclusions
In this paper, we focused on the stability of the lattice orientation for velocity gradient driven processes. To simplify the problem, we considered only rigid visco-plastic models, without hardening/softening effects. We formulated the general problem of lattice orientation stability analysis explicitly in a mathematical framework where tools such as linear stability could be used. This analysis allowed us to select the attractors from among the stationary lattice orientations and hence to predict the final texture of the poly/monocrystal for a given velocity gradient driven process.
However, we found that applying this stability analysis to a specific 3-D crystal (FCC, HCP, etc.) is very complex due to two main difficulties: the analytic description of the slip rates as functions of lattice orientation and the non-linearity of the problem. The detailed analysis of the 3-D case, which is beyond the scope of our present paper, could be considered in future work. For this paper, we restricted ourselves to the simplified 2D model with three slip systems. For this model, we developed analytic formulas of the slip rates as functions of the lattice orientation and the strain rate. We used this key result in our linear stability analysis to characterize the stable stationary orientations and to compute their attractors. Since the attractors are effective (i.e., the distance between the lattice orientation and the attractor is small) only for slip strains larger than 50%, this stability analysis, which concerns large deformations, is not so useful for small strains.
Figure 15: The computed distribution of the accumulated plastic slip $\gamma^{s}_{a}$ ($\dot{\gamma}_{a}^{s}=|\dot{\gamma}^{s}|$) at the end of the process $t=T$ ($\epsilon^{eng}=0.2$) for slip systems: $s=1$ (left), $s=2$ (center), and $s=3$ (right).
To illustrate our theory, we performed numerical simulations for three specific problems. In all cases, we gave only the boundary conditions with respect to the driven velocity field, but the computed velocity field differed from the expected one. However, we could deduce the overall final orientation of the lattice from the a priori attractor distribution predicted by our stability analysis.
In the first case, we considered a homogeneous driven velocity gradient field with non-homogeneous initial orientations, modeling the grains of a polycrystal. Since we initialized all grains with orientations belonging to one of two attractors, our stability analysis yielded the expected final texture of the polycrystal. We found the computed orientation to be in very good agreement with our predicted one, with a gap (in $L^{2}$ norm) of less than $4^{\circ}$. We also analyzed the situation when only one grain had an initial unstable orientation. We observed that this grain developed many shear bands, but globally it had a limited effect on the global orientation of the polycrystal. This comparison illustrated how subtle variations in initial crystallographic orientations can significantly affect the distribution of strain within individual grains.
In the second case, we studied a mono-crystal (homogeneous initial orientation) with a non-homogeneous driven velocity gradient: void growth under radial loading. Our numerical simulations showed that, in accordance with our stability analysis, the orientation at large strains exhibited a discontinuity at the angles $\varphi=\pi/6+k\pi/3$, as the attractor does. We observed these discontinuities near the void, where the deformations were large enough, and they vanished near the outer boundary, where the deformation was too small for the attractor to be effective. Very close to the void, at large strains, we found that the deformation process was much more complex: slip and buckling bands appeared and the strain rate differed significantly from the driven one. Therefore, we measured a larger difference between the lattice orientation and the estimated attractor very close to the void.
The third case we investigated was a slip band in uni-axial traction of a mono-crystal. Our stability analysis of the slip band predicted that the lattice orientation of the crystal in the band would rotate such that one slip system would become parallel with the slip band. For large strains, we confirmed that this slip system was the only active slip system in the slip band. However, we identified some situations for FCC crystal where there existed a small miss-orientation ($4^{\circ}$) between the slip system and the slip band. In our analysis, we supposed the slip band direction to be known and independent of the lattice rotation. But generally, we expect a complicated interplay between the sample shape, the boundary conditions, and the initial crystal orientation which generates the existence of one or several slip bands and their slopes.
We designed our numerical simulations to illustrate the assumption of two rigid blocks separated by one slip band with a high level of localized plastic strain (more than 200%) such that our stability analysis could be pertinent. As we expected, the orientation in the slip band changed during the slip to reach asymptotically the attractor with only one active slip system.
Finally, our theoretical stability analysis, illustrated by our numerical simulations, provides a comprehensive investigation of how initial crystallographic orientations, deformation mechanisms, and applied loading conditions affect the evolution of lattice orientations. Moreover, our theoretical knowledge of the attractors can greatly help predict crystal texture at large strains.
Acknowledgments. I. R. I. thanks Prof. Armand Beaudoin for interesting and useful discussions, which generated his interest in this subject. O. U. S. was supported by grants ANR-18-CE42-0017-03, ANR-19-CE08-0010-01, and ANR-20-CE91-0010.